OpenStudy (anonymous):

Sorry for being a pain :D Use the substitution method to solve the following system of equations. 4x − y = 3 7x − 9y = −2 (1, 1) (6, −3) (6, 1) (1, −3)

hero (hero):

Add y to both sides, then subtract 3 from both sides to isolate y in the first equation. Doing so gives: y = 4x - 3 The expression for y can be substituted into the second equation: 7x - 9y = -2 and after substitution we have 7x - 9(4x - 3) = -2 Now we simplify and solve for x: 7x - 36x + 27 = -2 (7 - 36)x + 27 = -2 -(36 - 7)x + 27 + 2 = 0 -(29)x + 29 = 0 29 = 29x 29/29 = x 1 = x Now plug the value for x back into the one of the original equations to get y.

OpenStudy (anonymous):

I don't know how to get the Y... that's why im asking for help..

OpenStudy (anonymous):

Do I do it the same way you got x?

hero (hero):

We isolated y earlier to get y = 4x - 3 If we substitute x = 1 into that equation, we get y.

OpenStudy (anonymous):

Oh.. so its D

OpenStudy (anonymous):

(1, -3)

hero (hero):

y = 4(1) - 3 y = 4 - 3 y = 1 Therefore, the point is (1,1)

OpenStudy (anonymous):

Im confused

hero (hero):

Re-read it all over again, but slowly.